The total energy of a particle, executing simple harmonic motion is________________?

Grade:11

5 Answers

Revanth - -
11 years ago
Dear shiva s

This is just a formula
total energy of a particle executing shm is TE = 1/2 M(ω)2 A2
= ALWAYS CONSTANT
LIKE IF YOU THINK MY ANSWER IS RIGHT
upender surepally
11 years ago
In simple harmonic motion when  a particle is displaced to a position from its mean position, then its kinetic energy gets converted into potential energy and vice-versa. Hence, total energy of a particle remains constant or the total energy in simple harmonic motion does not depend on displacement x.
Kanvi
6 years ago
Kinetic energy = 1/2mv^2 Where V =velocity =w^2(r^2 -y^2)Potential energy =1/2 Ky^2K=spring or force constant Total energy=1/2mw^2r^2Also m w^2 =K Thus expression is 1/2 K r^2
Rajneesh Singh
6 years ago
In SHM, acceleration =w^2 yAt extreme position y=a and whole energy is potential velocity,v = w √a^2-y^2At mean position y =0 , the whole energy is kineticKE at mean position=1/2m(velocity)^2Total energy = 1/2 m (wa)^2 = 1/2m^2a^2Total energy is directly proportional to (amplitude)^2
Kushagra Madhukar
3 years ago
Dear student,
Please find the attached solution to your problem.
 
In simple harmonic motion when  a particle is displaced to a position from its mean position, then its kinetic energy gets converted into potential energy and vice-versa. Hence, total energy of a particle remains constant or the total energy in simple harmonic motion does not depend on displacement x.
The formula for total energy of a particle executing shm is TE = 1/2 M(ω)2 A2
 
Hope it helps.
Thanks and regards,
Kushagra